# Asynchronous mass inversion enriched quantum anomalous Hall states in multilayer graphene

### v1

@@ Abstract
The low-energy bands near the $K$ valley for RPG without and with $\gamma_{3}$ are shown in Fig. S1 (b) (and (c)). From these figures, we can conclude that the $\gamma_{3}$-trigonal warping term induces further splitting of the central touching point. However, while this splitting transforms the quadratic band-touching at the $K$ point into a central Dirac cone and three “leg” Dirac cones located at and around the $K$ point, the momentum amplitude for the three “leg” Dirac cones remains very small due to the condition $v_{0}\gg v_{3}$. Consequently, the condition for mass inversion in the central Dirac cone, $\Delta=-m$, is very close to the condition for mass inversion in the “leg” Dirac cones, $D=\frac{2\lambda-1}{1+\lambda}m$, where $\lambda=\frac{v_{0}^{2}a_{0}^{2}k^{2}}{2\gamma_{1}^{2}}\ll 1$. This proximity results in the same phase diagram as in the case of $\gamma_{3}=0$ when the displacement field increases at a fixed staggered layer order.
SUPPLEMENTAL MATERIAL NOTE 2: BERNAL MULTILAYER GRAPHENE
.4 low-energy effective model for Bernal multilayer graphene
In this section, we derive the low-energy effective model for Bernal graphene with an even number of layers. The Hamiltonian for Bernal multilayer graphene (BMG) in the basis $\psi=(A_{1},B_{1},A_{2},B_{2},\dots,A_{N},B_{N})$ is written as:
@@
For $N$-layer BMG, as discussed in Ref.[71], we can decompose the Hamiltonian into a nearly block-diagonal form by using a unitary transformation. For $N\in Even$, the new basis is constructed by combining the old basis as follows:
@@
.5 The band structure and topological phase transition for the full model of the Bernal tetralayer graphene
In this section, we calculate the band structure for the full model of Bernal tetralayer graphene (Eq. (S11)) for the case where $N=4$, and investigate the topological phase transition that occurs as the displacement field increases in the presence of staggered layer orders (Eq. (S3)). From the results shown in Fig. S4, we can conclude that the QAH state with a total Chern number $C=6$ persists.

### v2

@@ Abstract
The low-energy bands near the $K$ valley for RPG without and with $\gamma_{3}$ are shown in Fig. S1 (b) (and (c)). From these figures, we can conclude that the $\gamma_{3}$-trigonal warping term induces further splitting of the central touching point. However, while this splitting transforms the quadratic band-touching at the $K$ point into a central Dirac cone and three satellite Dirac cones located at and around the $K$ point, the momentum amplitude for the three satellite Dirac cones remains very small due to the condition $v_{0}\gg v_{3}$. Consequently, the condition for mass inversion in the central Dirac cone, $\Delta=-m$, is very close to the condition for mass inversion in the satellite Dirac cones, $D=\frac{2\epsilon-1}{1+\epsilon}m$, where $\epsilon=\frac{v_{0}^{2}a_{0}^{2}k^{2}}{2\gamma_{1}^{2}}\ll 1$. This proximity results in the same phase diagram as in the case of $\gamma_{3}=0$ when the displacement field increases at a fixed staggered layer order.
.4 I-D. The entire phase diagram under increasing displacement field for RPG
Figure S3: (a). The evolution of the band structure with increasing interlayer potential at $K$ valley, in the presence of the staggered sublattice order given in Eq. (S11). The parameters are $m=m_{1}=-15~{}meV$, $\gamma_{2}=-20~{}meV$, $\gamma_{0}=3160~{}meV$, and $\gamma_{1}=390~{}meV$. (b). The evolution of the band structure with increasing interlayer potential in the case where $\gamma_{2}=0$. All other parameters are the same as in (a).
Using the staggered layer order defined in Eq. (S3), we can explain the topological phase transitions from the $C=0$ state to the $C=-3$ QAH state, and then to the $C=-5$ QAH state. However, as the displacement field continues to increase, the transition from the $C=-5$ to the $C=-3$ QAH state, and then to the $C=0$ state, remains unexplained Han et al. (2024a). To account for this sequence of topological phase transitions, it is necessary to introduce an order parameter that first induces mass inversion at the central point. One natural choice is the staggered sublattice order, which describes a uniform charge density across each layer but with opposite charge polarizations between the $A$ and $B$ sublattices within a single layer. This order also satisfies the charge neutrality condition. The specific form of the Hamiltonian for this order parameter in the basis $\psi=(A_{1},B_{1},A_{2},B_{2},\dots,A_{5},B_{5})$ is written as:
$H^{\prime}_{m}=m~{}Diag(\sigma_{z},\sigma_{z},...,\sigma_{z}),$ (S11)
where $\sigma_{z}$ is the Pauli matrix of the sublattice degree of freedom on each layer. By performing perturbation theory, the effective mass for the low-energy effective model is derived as:
$M^{\prime}_{5,\text{eff}}(k)=(m+\Delta)+\frac{(2m+\Delta)v_{0}^{2}k^{2}}{2\gamma_{1}^{2}}.$ (S12)
For $k=0$ point, the effective mass changes sign at $\Delta=\Delta^{\prime}_{1}=-m$ (assuming $\Delta>0$ and $m<0$). However, for $k\neq 0$ points, unlike the case of the staggered layer order, the effective mass changes sign at $\Delta=\Delta^{\prime}_{2}=-m\frac{1+2\epsilon}{1+\epsilon}$ which is larger than $\Delta^{\prime}_{1}=-m$ due to $\epsilon=\frac{v_{0}^{2}a_{0}^{2}k^{2}}{2\gamma_{1}^{2}}>0$. Thus, as the displacement field increases, the effective mass at the central point of the $K$ (or $K^{\prime}$) valley changes sign first at $\Delta=\Delta_{1}^{\prime}=-m$, followed by the sign change of the effective mass of the three satellite Dirac cones at $\Delta=\Delta^{\prime}_{2}>\Delta_{1}^{\prime}$. In the presence of the staggered sublattice order, the evolution of the band structure at the $K$ valley with increasing interlayer potential is shown in Fig. S3 (a) and (b), for cases with and without trigonal warping, respectively.
We then assume that the staggered orders for the four flavors $K\uparrow$, $K\downarrow$, $K^{{}^{\prime}}\uparrow$, and $K^{{}^{\prime}}\downarrow$ are given as $H^{\prime}_{m_{1}}$, $-H^{\prime}_{m_{1}}$, $H_{m_{2}}$, and $-H_{m_{2}}$, respectively, with $0>m_{2}>m_{1}$. Under this assumption, the order in the $K$ valley corresponds to staggered sublattice order, while in the $K^{\prime}$ valley, it corresponds to staggered layer order. When the interlayer potential satisfies $0<\Delta<-m_{2}$, as $\Delta$ increases, mass inversions occur at the three satellite Dirac cones for $K^{\prime}\uparrow$ flavor, causing the Chern number of RPG to change from $C=0$ to $C=-3$. Then, when $\Delta=-m_{2}$, the effective mass at the central band-touching point for electrons with flavor $K^{\prime}\uparrow$ changes sign, leading to a further change in the Chern number from $C=-3$ to $C=-5$. When the displacement field satisfies $-m_{2}<\Delta<-m_{1}$, since $m_{2}>m_{1}$, the Chern number for RPG remains $C=-5$. As the displacement field continues to increase, for electrons with flavor $K\uparrow$, the effective mass of the central band-touching point changes sign at $\Delta=-m_{1}$, leading to a topological phase transition from the QAH state with $C=-5$ to $C=-3$. With further increases in the displacement field, the effective masses of the three satellite Dirac cones with flavor $K\uparrow$ change sign, resulting in a topological phase transition from the QAH state with $C=-3$ to the state with $C=0$. Consequently, we deduce that as the displacement field increases, the Chern number for RPG evolves as $0\rightarrow-3\rightarrow-5\rightarrow-3\rightarrow 0$, which is consistent with experimental observations Han et al. (2024a). Moreover, within this mechanism, we can derive that at $\Delta=0$, the $C=0$ state in RPG corresponds to a state where both the spin and valley Chern numbers are 0. When the displacement field becomes sufficiently large, although the total Chern number of RPG remains 0, the system transitions into a quantum valley Hall state with a valley Chern number of 5.
SUPPLEMENTAL MATERIAL NOTE II: BERNAL MULTILAYER GRAPHENE
.5 II-A. low-energy effective model for Bernal multilayer graphene
In this section, we derive the low-energy effective model for Bernal multilayer graphene (BMG). For Bernal $N$-layer graphene, if $N$ is even, the Hamiltonian can be decomposed into $N/2$ mutually coupled bilayer graphene systems Koshino and McCann (2011). We can then project the Hamiltonian for each bilayer graphene onto the basis which is composed of non-dimerized “atoms” to obtain an $N$-dimensional low-energy effective model. If $N$ is odd, the low-energy effective model for BMG consists of $N/2$ mutually coupled bilayer graphene systems along with an additional monolayer graphene that couples to these bilayer graphenes. The presence of this additional monolayer graphene results in a metallic state, even when staggered layer orders are considered Grushina et al. (2015), thereby preventing the formation of a QAH state. As a result, this paper focuses on even-layer BMGs. The full Hamiltonian for $N$-layer BMG in the basis $\psi=(A_{1},B_{1},A_{2},B_{2},\dots,A_{N},B_{N})$ is written as:
@@
For $N$-layer BMG, as discussed in Ref.Koshino and McCann (2011), we can decompose the Hamiltonian into a nearly block-diagonal form by using a unitary transformation. For $N\in Even$, the new basis is constructed by combining the old basis as follows:
@@
.6 II-B. The band structure and topological phase transition for the full model of the Bernal tetralayer graphene
In this section, we calculate the band structure for the full model of Bernal tetralayer graphene (Eq. (S13) with the conduction $N=4$), and investigate the topological phase transition that occurs as the displacement field increases in the presence of staggered layer orders (Eq. (S3)). From the results shown in Fig. S5, while the band structure and topological phase transition points differ from those in the low-energy effective model, the QAH state with a total Chern number of $C=6$ persists.
Figure S6: (a), (b), (c). The band structure for the full model of Bernal tetralayer graphene under the increasing displacement field in the presence of the staggered layer orders $m=6~{}meV$ in $K$ valley. (d), (e), (f). The band structure for the full model of Bernal tetralayer graphene under the increasing displacement field in the presence of the staggered layer orders $m=-6~{}meV$ in $K^{\prime}$ valley. The Chern number for each valence band is indicated near the respective band. The hopping parameters are: $v_{0}=\frac{\sqrt{3}}{2}\times 3000~{}meV$, $\gamma_{1}=400~{}meV$, $\gamma_{2}=-20~{}meV$, $v_{3}=\frac{\sqrt{3}}{2}\times 300~{}meV$, $v_{4}=\frac{\sqrt{3}}{2}\times 40~{}meV$, and $\gamma_{5}=40~{}meV$.
Additionally, the topological phases and band structure in the full model of BTG are influenced by the amplitude of the staggered layer order. In Fig. S6, compared to the case in Fig. S5, we choose staggered layer orders for the $K$ and $K^{\prime}$ valleys with smaller amplitudes ($|m|=6~{}meV$). It turns out that the QAH state with Chern number $C=4$ disappears, while the QAH state with Chern number $C=6$ persists, which is consistent with the experimental observations Cho et al. (2025). In the case of $|m|=6~{}meV$, the next-nearest interlayer hoppings $\gamma_{2}$ and $\gamma_{5}$ significantly affect the effective masses of the first and second bilayer graphene, preventing the mass inversion at the central Dirac cone of the first bilayer graphene. The topological phase transition in the $K$ valley is driven by the mass inversion of the satellite Dirac cones belonging to the bands of the first bilayer graphene, resulting in a change of the Chern number by $3$.